OpenStudy (anonymous):

How do i differentiate x-2/x^2

OpenStudy (perl):

do you mean (x-2)/x^2 or x - 2/x^2

OpenStudy (anonymous):

(x-2)/x^2

OpenStudy (gorv):

\[\frac{ d \frac{ u }{ v}}{ dx }=\frac{ v*\frac{ du }{ dx } - u*\frac{ dv }{ dx }}{ v^2 }\]

OpenStudy (gorv):

u=x-2 v=x^2

OpenStudy (gorv):

can u do it now ??

OpenStudy (anonymous):

No sorry. I get to the stage where i have simplified it to be \[\frac{ 1 }{ x } - \frac{ 2 }{ x^2 }\] But i don't know how to get the rational expression into a negative power.

OpenStudy (gorv):

u got the wrong ans

OpenStudy (gorv):

u=x-2 and v= x^2 find du/dx and dv/dx and put in the formula i gave earlier

OpenStudy (anonymous):

I have converted the \[\frac{ 1 }{ x }\] to \[x^-1\] but i don't know how to convert \[\frac{ 2 }{ x^2 }\] . Can you please show me how you can convert it... I know how to do the rest....

ganeshie8 (ganeshie8):

\[\large \dfrac{1}{a^n} = a^{-n}\]

ganeshie8 (ganeshie8):

\[\large \dfrac{2}{x^2} = 2 \times \dfrac{1}{x^2} = 2 \times x^{-2}\]

OpenStudy (anonymous):

is there a reasons why you times by 2 if the original question is divided by 2

ganeshie8 (ganeshie8):

multiply \(\large a\) and \(\large \dfrac{1}{b}\) what do you get ?

OpenStudy (anonymous):

a/ab

OpenStudy (anonymous):

\[\frac{ a }{ab }\]